OpenStudy (anonymous):

A pineapple was found to have two sets of opposing spirals formed by its leaves. The clockwise spiral has x leaves. The number of leaves in the counterclockwise direction is 13 more than that in the clockwise direction. These leaf numbers are consecutive numbers in the Fibonacci sequence. Which equation is true for the number of leaves in the two spirals? Answer The quantity x plus 13 divided by x equals 1.6 X divided by 13 equals 1.6 The quantity x plus 13 divided by 13 equals 1.6 13 divided by x equals 1.6

OpenStudy (anonymous):

I know it is the Fibonacci sequence thing

OpenStudy (ash2326):

The clockwise spiral has x leaves. The number of leaves in the counterclockwise direction x+13 fibonacci series is 0,1,1,2,3,5,8,13,21,34 the second term is generated by the sum of the two previous terms 0,1,0+1,1+1,1+2 and so on since x and x+13 are also terms of fibonacci series x+13 is =x + 13 where x is term before x+13 13 is term before x so from the series we can see x=21 x+13=34 so no. of clockwise leaves=21 counter clockwise leaves=34 x+13=34 34/21=1.6 so first option

OpenStudy (anonymous):

Some say Fib series goes like 1,1,2,3,5,8,13,21,34

OpenStudy (ash2326):

yeah both are coreect , it may start friom zero or 1

OpenStudy (anonymous):

and the answer is golden ratio. ( http://en.wikipedia.org/wiki/Golden_ratio) !!